Estimated demand function for processed pork is Q=171-20p+20pb+3pc+2y. Using the estimated demand function for processed pork, show how the quantity demanded at a given price changes as price of Beef (pb), decreases by $ 0.95 a year?


1
Expert's answer
2022-03-23T18:34:07-0400

Q=171-20p+20pb+3pc+2y ,where pb is the price for beef

lets assume initially the price of beef was pb=$2,then

Q=211-20p+3pc+2y, you realize that the autonomous quantity of pork demanded has increased by 40 units.

:.now lets say this prices decrease by $0.95,then 2-0.95=1.05 and the function

Q=192-20p+3pc+2y,the autonomous quantity demanded for pork has decreased by 19 units ,this shows that pork and beef are substitute commondities and a decrease of price for one will lead to decrease of quantity demanded for the other.